# Flatness-based Finite-Horizon Multi-UAV Formation Trajectory Planning and Directionally Aware Collision Avoidance Tracking

### v1

@@ Abstract
Collision-free optimal formation control of unmanned aerial vehicle (UAV) teams is challenging. The state-of-the-art optimal control approaches often rely on numerical methods sensitive to initial guesses. This paper presents an innovative collision-free finite-time formation control scheme for multiple UAVs leveraging the differential flatness of the UAV dynamics, eliminating the need for numerical methods. We formulate a finite-time optimal control problem to plan a formation trajectory for feasible initial states. This formation trajectory planning optimal control problem involves a collective performance index to meet the formation requirements of achieving relative positions and velocity consensus. It is solved by applying Pontryagin’s principle. Subsequently, a collision-constrained regulating problem is addressed to ensure collision-free tracking of the planned formation trajectory. The tracking problem incorporates a directionally aware collision avoidance strategy that prioritizes avoiding UAVs in the forward path and relative approach. It assigns lower priority to those on the sides with an oblique relative approach and disregards UAVs behind and not in the relative approach. The simulation results for a four-UAV team (re)formation problem confirm the efficacy of the proposed control scheme.
@@ 1 Introduction
Formation control of unmanned aerial vehicles (UAVs) has received growing interest in recent decades with a broad range of applications in exploring hazardous terrain [1], transporting loads [2], performing perilous inspection tasks [3], surveying structures of historical significance [4], and more. In formation trajectory planning, UAVs plan optimal or sub-optimal collision-free trajectories between their initial and terminal states [5, 6]. In formation trajectory tracking, all UAVs converge to the pre-planned formation trajectories [7, 8, 9].
The nonlinear dynamics of UAVs, inter-UAV collision-free coordination, and UAV-obstacle collision avoidance pose additional challenges to formation trajectory planning and tracking. Finite-horizon formation control, achievable by optimizing a collective performance index consisting of UAV states and control inputs within a finite time frame, holds practical significance [10]. The time frame is constrained when UAV groups plan to create a formation to execute maneuvers, reach destinations, optimize resource utilization (e.g., energy), and, most importantly, ensure adherence to safety constraints. Consequently, collision-free finite-time formation trajectory planning and tracking are a frontier of ongoing research.
Optimal trajectory planning for UAV systems is difficult, and solutions often rely on numerical methods that can be sensitive to initial guesses [11]. Instead, most studies on formation trajectory planning focus on either interpolating or approximating the optimal trajectories between the initial and terminal states [12]. In [12, 5], algorithms are developed for generating collision-free finite-time trajectories for multi-UAV systems, utilizing Bézier curves [12], high-order piecewise polynomials [5], and piecewise polynomial trajectory generation for UAVs recursively [6] to parameterize trajectories between the initial and terminal states. However, the literature on UAV formation control lacks approaches for collision-avoidance finite-horizon formation trajectory planning and tracking of UAVs solely based on their initial states.
This paper leverages the inherent differential flatness property of quadrotor-type UAV dynamics to introduce a finite-horizon formation trajectory planning and tracking control scheme solely based on UAVs’ initial states. The finite-horizon planning trajectories in the flat coordinates are derived by applying necessary and sufficient conditions for optimality using Pontryagin’s principle. The trajectory tracking control system is optimal when all UAVs operate within their safety region and becomes sub-optimal when any two UAVs enter their collision reaction region. The paper’s main contribution is an inter-UAV collision-free optimal finite-horizon formation control scheme for formation trajectory planning and tracking using solely the initial state of UAVs. The proposed approach
@@
    is consensus-based and, unlike [13, 14, 15, 16], does not rely on a (virtual or actual) leader, which typically serves as the external reference for the formation, thereby enhancing robustness and scalability by eliminating potential single-point failures,
@@
    unlike [12, 5, 6], solely uses the initial states of the UAVs to plan optimal trajectories, which reduces computational complexity and eliminates the need for external references or predefined waypoints,
@@
    unlike [16], employs a finite horizon, which enables real-time adaptability to changing tasks and dynamic environments with finite time convergence, and
@@
    finally, introduces a directionally aware collision avoidance strategy that prioritizes collision avoidance both in the forward path and the relative approach for each UAV. This results in an adaptive and efficient behavior that reduces or prevents unnecessary maneuvering toward UAVs on the sides or behind, as well as those not on a relative approach trajectory, which are less likely to pose an immediate threat or pose no threat. This aspect is not addressed in existing works such as [8, 17].
The rest of this paper is organized as follows. Section 2 provides preliminaries. The collision-free finite-time formation control is defined in Section 3. Section 4 recasts the problem definition as optimal control problems where their solution is investigated. The directionally aware collision avoidance strategy is introduced in Section 5. Section 6 illustrates the simulation results for a four-UAV team (re)formation. Section 7 concludes the paper.
@@ 2 Preliminaries
A directed graph is a pair $\mathcal{G}(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ is a finite set of vertices/nodes and $\mathcal{E}\subseteq\{(i,j):i,j\in\mathcal{V}\}$ is a set of directed edges/arcs. Each edge $(i,j)\in\mathcal{E}$ represents a directional information flow and is assigned with a positive scalar weight $\mu_{ij}>0$. An undirected graph is similarly defined with its edges being undirected, indicating bidirectional information flow. Graph $\mathcal{G}$ is connected meaning that for every pair of vertices $(i,j)\in\mathcal{V}\times\mathcal{V}$, from $i$ to $j$ for all $j\in\mathcal{V},j\neq i$, there exists a path of (undirected) edges from $\mathcal{E}$.
@@
$\mathbf{L}=\mathbf{D}\mathbf{W}\mathbf{D}^{\top}\in\mathbb{R}^{N}.$
Matrix $\mathbf{D}\in\mathit{\mathbb{R}}^{\mathcal{|V|}\times\mathcal{|E|}}$ is the incidence matrix of $\mathcal{G}$ where $\mathbf{D}$’s $ij$th element is $1$ if the node $i$ is the tail of the edge $j$, $-1$ if the node $i$ is the head, and 00, otherwise. For an undirected graph, the designation of the tail or head is arbitrary. Matrix $\mathbf{W}=\mathrm{diag}(\cdots,\mu_{ij},\cdots)~{}\forall(i,j)\in\mathcal{E}\in\mathit{\mathbb{R}}^{|\mathcal{E}|}$. Matrix $\mathbf{L}$ is symmetric ($\mathbf{L}=\mathbf{L}^{\top}$), positive semidefinite (PSD) ($\mathbf{L}\succeq\mathbf{0}$), and satisfies the sum-of-squares property
@@
Let $\mathbf{Y}=[y_{ij}]\in\mathit{\mathbb{R}}^{p\times q}$. Thus,
$\mathbf{Y}\otimes\mathbf{I}_{m}=\begin{bmatrix}y_{11}\mathbf{I}_{m}&\cdots&y_{1q}\mathbf{I}_{m}\\ \vdots&\ddots&\vdots\\ y_{p1}\mathbf{I}_{m}&\cdots&y_{pq}\mathbf{I}_{m}\end{bmatrix}\in\mathit{\mathbb{R}}^{mp\times mq}.$
@@ 2.2 UAV model
Figure 1: Quadrotor configuration. $\mathcal{W}$ and $\mathcal{B}$ denote the world frame and body frame, respectively. The roll, pitch, and yaw angles of the vehicle are denoted by $\phi_{i}(t)$, $\theta_{i}(t)$, and $\psi_{i}(t)$, respectively.
@@ 3 Problem definition
Consider a multi-UAV team of $N$ quadrotors indexed by $\mathcal{V}=\{1,\cdots,N\}$. Each UAV $i\in\mathcal{V}$ is described by the dynamics (2)--(3). Two time-invariant connected directed graphs are defined: the formation graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ and the communication graph $\mathcal{G}_{c}(\mathcal{V},\mathcal{E}_{c})$. The formation graph specifies the desired spatial formation shape, while the communication graph outlines the inter-UAV communications.
@@
The formation graph is connected.
@@
The communication graph is complete to account for all inter-UAV collisions (thus, $\mathcal{G}_{f}\subseteq\mathcal{G}_{c}$).
The collision-free formation control problem consists of steering all UAVs from their initial state to a desired formation while avoiding inter-UAV collisions. Formation shape is determined by the offset vectors $\mathbf{d}_{ij}\in\mathbb{R}^{3}$ for all UAV pairs $(i,j)\in\mathcal{E}_{f}$. Each UAV $i\in\mathcal{V}$ has an avoidance region radius $r_{i}$, meaning the minimum safe distance between two UAVs $i$ and $j$ is $r_{i}+r_{j}$. Let $R_{i}$ denote the radius of the collision reaction region of each UAV $i\in\mathcal{V}$ and $R_{i}>r_{i}$ (see Fig. 2).
@@
The multi-UAV formation control problem is defined as follows.
@@
(a) Consider a multi-UAV system on the directed connected graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ with UAV dynamics (2)--(3). The multi-UAV formation control problem is to determine control strategies that drive all UAVs from their initial state to a desired formation specified by $\mathbf{d}_{ij},\forall(i,j)\in\mathcal{E}_{f}$ in the state space in a finite time horizon $t_{f}>0$. (b) All UAVs must avoid mutual collisions, i.e., $\Lambda(\mathbf{p})$ and $\Phi(\mathbf{p})$ must be $\emptyset$ for all $t\in(0,t_{f})$.
The formation control objective is to ensure
@@
The UAV dynamics, formation control objective, and collision avoidance constraint described by (2)--(5) can be recast as a constrained optimal control problem. This allows for investigating optimal formation control strategies for Problem 1. However, finding a solution is problematic due to the nonlinear UAV dynamics. Numerical methods can be sensitive to initialization [11] and unstable, often relegating them to a lower priority[19, 20]. Moreover, the quadrotor system is inherently underactuated, with six degrees of freedom but only four control inputs. This characteristic poses an additional challenge in finding a solution that adheres to the system’s underactuation dynamics.
@@ 4 Collision-Free Trajectory Planning and Tracking
Leveraging the differential flatness property of the UAV dynamics provides a fully actuated system with four flat outputs [18, 21, 22, 14]. By applying a diffeomorphism (see Definition 1 from [23]), Problem 1 is transformed into a flat space where it becomes easier to handle due to the linear UAV dynamics. The simplified problem, denoted as Problem 2, is recast and addressed as optimal control trajectory planning and tracking problems in flat space. The flat space control strategies are then transformed back to the original coordinate system using the inverse of the diffeomorphism. This approach is summarized in the control scheme diagram in Fig. 3.
@@
(a) Consider a multi-UAV system on the directed connected graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ in flat space with dynamics (6). The multi-UAV formation control problem is to determine control strategies for the UAV team that drive all UAVs from their initial state to a desired formation specified by $\mathbf{d}_{ij},\forall(i,j)\in\mathcal{E}_{f}$ in the flat state space in a finite time horizon $t_{f}$. (b) Additionally, $\Lambda(\mathbf{p})$ and $\Phi(\mathbf{p})$ must be $\emptyset$ for all $t\in(0,t_{f})$.
@@
The convexity of the objective functional (7) and dynamics (6) makes Pontryagin’s principle both necessary and sufficient for optimality. Applying these conditions, the optimal formation trajectory planning signal for any initial state of the UAV team is presented as follows.
@@
Consider the multi-UAV formation control problem in flat coordinates defined by (6) and (7). The optimal formation trajectory for any initial state $\mathbf{r}_{0}$ is given by

### v2

@@ Abstract
Optimal collision-free formation control of the unmanned aerial vehicle (UAV) is a challenge. The state-of-the-art optimal control approaches often rely on numerical methods sensitive to initial guesses. This paper presents an innovative collision-free finite-time formation control scheme for multiple UAVs leveraging the differential flatness of the UAV dynamics, eliminating the need for numerical methods. We formulate a finite-time optimal control problem to plan a formation trajectory for feasible initial states. This optimal control problem in formation trajectory planning involves a collective performance index to meet the formation requirements to achieve relative positions and velocity consensus. It is solved by applying Pontryagin’s principle. Subsequently, a collision-constrained regulating problem is addressed to ensure collision-free tracking of the planned formation trajectory. The tracking problem incorporates a directionally aware collision avoidance strategy that prioritizes avoiding UAVs in the forward path and relative approach. It assigns lower priority to those on the sides with an oblique relative approach, disregarding UAVs behind and not in the relative approach. The high-fidelity simulation results validate the effectiveness of the proposed control scheme.
@@ 1 Introduction
Unmanned aerial vehicle (UAV) formation control has gained significant attention due to its broad applications, such as exploring hazardous environments [1], transporting loads [2], conducting critical inspections [3], and surveying historical structures [4]. A fundamental aspect of UAV formation control involves trajectory planning and trajectory tracking. In formation trajectory planning, UAVs plan optimal or sub-optimal collision-free trajectories between their initial and terminal states [5, 6]. Formation trajectory tracking ensures that UAVs converge to these pre-planned formation trajectories [7, 8, 9]. Formation trajectory planning and tracking present significant challenges due to the nonlinear dynamics of UAVs, inter-UAV collision avoidance, and dynamic environmental constraints. For quadrotor UAVs, their underactuated nature further complicates control and trajectory planning. These challenges become more pronounced in time-sensitive operations, where UAV groups must rapidly form formations, execute tasks, and ensure collision-free motion within finite time horizons [10].
A quadrotor UAV, with six degrees of freedom and only four control inputs, is inherently underactuated. Differential flatness resolves underactuation by representing states and inputs as functions of flat outputs (e.g., position and yaw) and their derivatives. This simplifies trajectory planning and control [11, 12, 13]. The application of differential flatness theory to quadrotor formation control has garnered research attention. The study in [14] has tackled the leader-follower formation control for quadrotors using differential flatness theory and proposed an observer-based controller to reconstruct the leader’s states and reject disturbances. In [15], a leader-follower formation control strategy is presented for a team of quadrotors, integrating trajectory planning (using the flatness-based trajectory generation algorithm in [11]), distributed consensus, and trajectory tracking. In [16], a leader-follower control strategy is developed for UAV formation in GPS-denied environments, incorporating feedback linearization, differential flatness, and linear-quadratic regulator (LQR) techniques. In [17], a virtual structure-based approach is introduced to generate formation trajectories in a quadrotor swarm, where quadrotors utilize differential flatness for control. The virtual structure is integrated with multiple layered potential fields to ensure collision avoidance.
Networked UAVs can exchange information with neighbors, enabling consensus algorithms to facilitate collaboration. These algorithms can be integrated with leader-follower [18], behavioral, or virtual structure methods for trajectory tracking or employed independently for trajectory planning and tracking in team formation control [19]. Finite-horizon formation control unifies consensus-based coordination with an optimal control framework by optimizing a collective performance index of UAV states and control inputs within a constrained time frame [10]. Some research applies discrete-time optimal control with fixed-time performance guarantees, such as the method for the seeker stabilized platform (SSP) in [20], which uses iterative updates and dynamic programming. In [21], a framework is proposed that uses fixed-time performance functions and adaptive neural approximations to efficiently handle unknown nonlinearities and directions, avoiding the reliance on Nussbaum-type functions commonly used in existing strategies for prescribed performance control of discrete-time non-affine systems. In [22], a discrete-time fuzzy-neural intelligent control approach with fixed-time prescribed performance is introduced, addressing the limitations of existing methods that struggle with time-varying sampling intervals. However, these approaches can be computationally intensive due to time-stepping algorithms [20], adaptive neural approximations [21], and fuzzy-neural design [22]. In contrast, our method uses continuous-time finite-horizon optimal control and leverages the differential flatness of UAV dynamics to enable efficient real-time coordination via linear-quadratic optimization.
Existing finite-horizon methods exhibit limitations due to their dependence on numerical optimization, leader-follower hierarchies, or pre-planned trajectory interpolations, each of which introduces practical constraints. Numerical methods for trajectory optimization require well-chosen initial guesses to converge to feasible solutions [23]. In the leader-follower approach [18], the leader serves as a single point of failure, the behavioral approach lacks a formalized control framework, and the virtual structure approach suffers from limited adaptability [14, 15, 17]. Trajectory interpolations using Bézier curves, high-order piecewise polynomials, or recursive constructions [24, 5, 6] require predefined initial and terminal states, and often waypoints. Existing collision-avoidance strategies often consider all neighboring UAVs equally without prioritizing evasive actions based on their relative motion [8, 25]. This can lead to inefficient maneuvers and unnecessary trajectory deviations. Although infinite-horizon approaches provide stability guarantees [26], they lack responsiveness to time-sensitive formation control tasks.
This paper addresses the above challenges by leveraging the differential flatness property of quadrotor UAV dynamics to develop a finite-horizon formation trajectory planning and tracking control scheme based solely on the initial states of the UAVs. Specifically, the proposed approach involves the following features:
@@
    Consensus-Based Formation Control: Unlike most prior works that rely on a virtual or actual leader as an external reference for formation control [27, 14, 28, 26], the proposed approach is entirely consensus-based. This improves robustness and scalability by eliminating potential single-point failures associated with leader-based strategies. Formation direction can be adjusted by incorporating the coordinates of the target area as a deviation term in the performance index, allowing the UAV group to maintain formation while approaching the target without relying on a designated leader.
@@
    Optimal Trajectory Planning Using Only Initial State Knowledge: The proposed framework optimally plans UAV trajectories using only their initial states by applying the necessary and sufficient conditions from the Pontryagin principle. This significantly reduces computational complexity, as trajectory planning and generation are performed only once.
@@
    Finite Horizon Real-Time Adaptability: Unlike infinite-horizon formulations [26], our approach employs a finite-horizon strategy, optimizing UAV motion within a limited time window to ensure timely responsiveness, adaptability to dynamic tasks and environments, and guaranteed finite-time convergence.
@@
    Directionally Aware Collision Avoidance: The proposed method prioritizes evasive actions based on relative motion, emphasizing avoidance maneuvers on the forward path and along the approach trajectory, thus reducing unnecessary deviations and ensuring efficient collision-free multi-UAV operations.
The main contribution of this paper is a unified framework for UAV formation control that integrates the features discussed above. It ensures collision-free trajectory planning and tracking while maintaining scalability, computational efficiency, and real-time adaptability. Furthermore, the approach leverages the linear-quadratic paradigm from optimal control, which is known for its analytic tractability and computational efficiency [29].
The rest of this paper is organized as follows. Section 2 provides preliminaries. The collision-free finite-time formation control is defined in Section 3. Section 4 recasts the problem definition as optimal control problems, where their solutions are investigated. The directionally aware collision avoidance strategy is introduced in Section 5. Section 6 presents illustrative examples and simulation results for the (re)formation of a four-UAV and seven-UAV team. Section 7 concludes the paper.
@@ 2 Preliminaries
A directed graph is a pair $\mathcal{G}(\mathcal{V},\mathcal{E})$ where $\mathcal{V}=\{1,\cdots,N\}$ is a finite set of vertices/nodes and $\mathcal{E}\subseteq\{(i,j):i,j\in\mathcal{V}\}$ is a set of directed edges/arcs. Each edge $(i,j)\in\mathcal{E}$ represents a directional information flow and is assigned a positive scalar weight $\mu_{ij}>0$. A graph $\mathcal{G}$ is connected if, for every pair of distinct vertices $(i,j)\in\mathcal{V}\times\mathcal{V}$ with $i\neq j$, there exists a path of undirected edges in $\mathcal{E}$ that connects $i$ to $j$.
@@
$\mathbf{L}=\mathbf{D}\mathbf{W}\mathbf{D}^{\top}\in\mathbb{R}^{N},$
where $\mathbf{D}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{E}|}$ is the incidence matrix of $\mathcal{G}$, with its $(i,j)$th element equal to $1$ if node $i$ is the tail of edge $j$, $-1$ if node $i$ is the head, and 00 otherwise. Additionally, $\mathbf{W}=\mathrm{diag}(\dots,\mu_{ij},\dots)$ for all $(i,j)\in\mathcal{E}$, where $\mathbf{W}\in\mathbb{R}^{|\mathcal{E}|}$. Matrix $\mathbf{L}$ is symmetric ($\mathbf{L}=\mathbf{L}^{\top}$), positive semidefinite (PSD) ($\mathbf{L}\succeq\mathbf{0}$), and satisfies the sum-of-squares property
@@
Let $\mathbf{Y}=[y_{ij}]\in\mathbb{R}^{p\times q}$. Then, the Kronecker product of $\mathbf{Y}$ and the identity matrix $\mathbf{I}_{m}$ is given by
$\mathbf{Y}\otimes\mathbf{I}_{m}=\begin{bmatrix}y_{11}\mathbf{I}_{m}&\cdots&y_{1q}\mathbf{I}_{m}\\ \vdots&\ddots&\vdots\\ y_{p1}\mathbf{I}_{m}&\cdots&y_{pq}\mathbf{I}_{m}\end{bmatrix}\in\mathbb{R}^{mp\times mq}.$
@@ 2.2 UAV model
Figure 1: Quadrotor configuration. $\mathcal{W}$ and $\mathcal{B}$ denote the world frame and body frame, respectively. The roll, pitch, and yaw angles are denoted by $\phi_{i}$, $\theta_{i}$, and $\psi_{i}$, respectively.
@@ 3 Problem definition
Consider a multi-UAV team consisting of $N$ quadrotors, indexed by $\mathcal{V}=\{1,\dots,N\}$. The dynamics of each UAV $i\in\mathcal{V}$ is governed by (2)–(3). Two time-invariant, connected, directed graphs are defined: the formation graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ and the communication or sensing graph $\mathcal{G}_{c}(\mathcal{V},\mathcal{E}_{c})$. The formation graph encodes the desired spatial relationships between UAVs, whereas the communication graph represents all potential sensing and communication links needed for collision avoidance.
@@
(Connectivity) The formation graph is directed, with its underlying undirected graph being connected and containing at least one globally reachable node, i.e., a node $i$ such that there is a directed path from $i$ to node $j$ for all $j\in\mathcal{V}$ with $j\neq i$ [30].
Assumption 1 implies the existence of a directed spanning tree with a root node $i$.
@@
The communication graph is a complete directed graph to account for all inter-UAV collisions (thus, $\mathcal{G}_{f}\subseteq\mathcal{G}_{c}$), where every pair of UAVs is connected by directed edges in both directions.
The formation control problem involves steering all UAVs from their initial state to a desired formation. Additionally, collisions between UAVs must be avoided. The formation shape is determined by the offset vectors $\mathbf{d}_{ij}\in\mathbb{R}^{3}$ for all UAV pairs $(i,j)\in\mathcal{E}_{f}$. Each UAV $i\in\mathcal{V}$ has an avoidance region radius $r_{i}$, meaning the minimum safe distance between two UAVs $i$ and $j$ is $r_{i}+r_{j}$. Additionally, let $R_{i}$ denote the radius of the collision reaction region of each UAV $i\in\mathcal{V}$ and $R_{i}>r_{i}$ (see Fig. 2).
@@
Assumption 3 ensures no collisions occur at the initial and final UAV positions. This guarantees the feasibility and safety of the formation configuration. By enforcing this condition, collision avoidance is required only during the transition. The multi-UAV formation control problem is defined as follows.
@@
(Multi-UAV Formation Control in State Space with Collision Avoidance) (a) Consider a multi-UAV system on the directed connected graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ with UAV dynamics (2)-(3). The multi-UAV formation control problem is to determine control strategies that drive all UAVs from their initial state to a desired formation specified by $\mathbf{d}_{ij},\forall(i,j)\in\mathcal{E}_{f}$ in the state space in a finite time horizon $t_{f}>0$. (b) All UAVs must avoid mutual collisions, i.e., $\Lambda(\mathbf{p})$ and $\Phi(\mathbf{p})$ must be $\emptyset$ for all $t\in(0,t_{f})$.
The consensus-based formation control objective is to ensure that
@@
The UAV dynamics, formation control objective, and collision avoidance constraint described by (2)-(5) can be recast as a constrained optimal control problem. This allows for investigating optimal formation control strategies for Problem 1. However, finding a solution is problematic due to the nonlinear UAV dynamics. Numerical methods can be sensitive to initialization [23] and unstable, often relegating them to a lower priority [31, 32]. Moreover, the quadrotor system is inherently underactuated, with six degrees of freedom but only four control inputs. This characteristic poses an additional challenge in finding a solution that adheres to the system’s underactuation dynamics.
@@ 4 Collision-Free Trajectory Planning and Tracking
Leveraging the differential flatness property of the UAV dynamics provides a fully actuated system with four flat outputs [11, 33, 14]. By applying a diffeomorphism (see Definition 1 from [34]), Problem 1 is transformed into a flat space where it becomes easier to handle due to the linear UAV dynamics. The simplified problem, denoted as Problem 2, is recast and addressed as optimal control trajectory planning and tracking problems in flat space. The flat space control strategies are then transformed back to the original coordinate system using the inverse of the diffeomorphism. This approach is summarized in the control scheme diagram in Fig. 3.
@@
(Multi-UAV Formation Control in Flat Space with Collision Avoidance) (a) Consider a multi-UAV system on the directed connected graph $\mathcal{G}_{f}(\mathcal{V},\mathcal{E}_{f})$ in flat space with dynamics (6). The multi-UAV formation control problem is to determine control strategies for the UAV team that drive all UAVs from their initial state to a desired formation specified by $\mathbf{d}_{ij},\forall(i,j)\in\mathcal{E}_{f}$ in the flat state space in a finite time horizon $t_{f}$. (b) Additionally, $\Lambda(\mathbf{p})$ and $\Phi(\mathbf{p})$ must be $\emptyset$ for all $t\in(0,t_{f})$.
@@
The objective functional (7) and the system dynamics (6) are both convex. Thus, Pontryagin’s principle is both a necessary and sufficient condition for optimality. Applying these conditions, the optimal formation trajectory planning signal for any initial state of the UAV team is presented as follows.
@@
Consider the multi-UAV formation control problem in flat coordinates defined by (6) and (7). The unique globally optimal control input and corresponding optimal formation trajectory for any initial state $\mathbf{r}_{0}$ are given by
$\mathbf{u}_{\mathbf{r}}(t)=-\mathbf{R}^{-1}\mathbf{B}^{\top}\mathbf{G}(t_{f}-t)\mathbf{H}^{-1}(t_{f})\mathbf{r}_{0},$ (8)
@@
$\mathbf{G}(t)=\begin{bmatrix}\mathbf{0}&\mathbf{I}\end{bmatrix}\exp{(-t\mathbf{M})}\begin{bmatrix}\mathbf{I}\\ \mathbf{Q}_{f}\end{bmatrix},$ (11)